# FourierKAN-GCF: Fourier Kolmogorov-Arnold Network - An Effective and Efficient Feature Transformation for Graph Collaborative Filtering

### v2

@@ Abstract.
Graph Collaborative Filtering (GCF) has achieved state-of-the-art performance for recommendation tasks. However, most GCF structures simplify the feature transformation and nonlinear operation during message passing in the graph convolution network (GCN). We revisit these two components and discover that a part of feature transformation and nonlinear operation during message passing in GCN can improve the representation of GCF, but increase the difficulty of training.
In this work, we propose a simple and effective graph-based recommendation model called FourierKAN-GCF. Specifically, it utilizes a novel Fourier Kolmogorov-Arnold Network (KAN) to replace the multilayer perceptron (MLP) as a part of the feature transformation during message passing in GCN, which improves the representation power of GCF and is easy to train. We further employ message dropout and node dropout strategies to improve the representation power and robustness of the model. Extensive experiments on two public datasets demonstrate the superiority of FourierKAN-GCF over most state-of-the-art methods. The implementation code is available at https://github.com/Jinfeng-Xu/FKAN-GCF.
@@
copyright: acmlicensedjournalyear: 2024doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation emai; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/18/06ccs: Information systems Recommender systems;
@@ 1. Introduction
Recommender system has been widely used to alleviate information overload on the web (Xu et al., 2024). The aim of recommender system is to predict whether a user will interact with an unobserved item. To this end, collaborative filtering (CF) addresses it by learning user preferences from similar users. Many works (He et al., 2020; Wang et al., 2019; Jiang et al., 2023) harnessing graph-based recommendation to capture latent user preferences from user-item interactions have gained notable successes. The main reason is that the interactive data between users and items are naturally graph-structured. For example, NGCF (Wang et al., 2019) utilizes whole standard GCN in recommender systems, which retains the feature transformation and nonlinear operation. However, LightGCN (He et al., 2020) states that both feature transformation and nonlinear operation are unnecessary for feature extraction in recommendation field and further proposes a light GCN. Most of the subsequent graph-based models (Zhou et al., 2023a; Wei et al., 2019; Yu et al., 2023) are based on LightGCN for further exploration.
@@
We provide an empirical analysis for NGCF and LightGCN in Section 2. Then, we state that feature transformation in NGCF can enhance the interaction representation and boost the performance of GCN, but the training process becomes more complex. In this work, we introduce a simple yet powerful graph-based recommendation model called FourierKAN-GCF. Particularly, FourierKAN-GCF incorporates a unique Fourier Kolmogorov-Arnold Network (KAN) in place of the traditional multilayer perceptron (MLP) within the feature transformation during message passing in GCNs. This substitution enhances the representational capabilities and training ease of the GCF. Additionally, we implement various dropout techniques to further boost the model’s representational strength and robustness. Extensive experiments on public datasets demonstrate the superiority of our FourierKAN-GCF over state-of-the-art methods.
@@ 2.1. NGCF Brief
NGCF retains feature transformation and nonlinear operation during the message passing in GCN, formally:
@@
where $\mathbf{e}_{u}^{(l)}$ and $\mathbf{e}_{i}^{(l)}$ represent the embedding of user $u$ and item $i$ after $l$ layers message propagation, respectively. $\mathcal{N}_{u}$ and $\mathcal{N}_{i}$ denote the interacted item set with $u$ and interacted user set with $i$, respectively. $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ are two trainable weight matrices to perform feature transformation in each layer. $\sigma$ is the nonlinear activation function. It is worth noting that $\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}$ and $\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}$ can be regarded as the aggregated representation from neighbors, while $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ and $\mathbf{W}_{2}(\mathbf{e}_{u}^{(l)}\odot\mathbf{e}_{i}^{(l)})$ can be regarded as the aggregated representation from interaction information. The final embeddings are calculated as:
(2) $\mathbf{\hat{e}}_{u}=\mathbf{e}_{u}^{(1)}||\mathbf{e}_{u}^{(2)}||...||\mathbf{e}_{u}^{(L)},\quad\mathbf{\hat{e}}_{i}=\mathbf{e}_{i}^{(1)}||\mathbf{e}_{i}^{(2)}||...||\mathbf{e}_{0}^{(L)},$
where ||||| | is the concatenation operation. NGCF only concatenates layer-$1$ to layer-$L$ and ignores ego layer-00, since ego layer-00 has already been considered in the first term $\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}$ and $\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}$ of message passing and propagation.
@@ 2.2. LightGCN Brief
LightGCN provides an analysis of the feature transformation and nonlinear operation in NGCF. It mainly provides four observations:
1. Removing the entire feature transformation including both $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ leads to consistent improvements over NGCF.
2. Removing only nonlinear operation $\sigma$ will lead to a small deterioration of performance.
3. Removing both entire feature transformation and nonlinear operation can improve performance significantly.
4. The deterioration of NGCF stems from the training difficulty rather than over-fitting.
To this end, LightGCN removes feature transformation and nonlinear operation during the message passing in GCN. Formally, LightGCN leverages the user-item graph to propagate embeddings as:
@@
where $\mathbf{e}_{u}^{(l)}$ and $\mathbf{e}_{i}^{(l)}$ represent the embedding of user $u$ and item $i$ after $l$ layers message propagation, respectively. $\mathcal{N}_{u}$ and $\mathcal{N}_{i}$ denote the interacted item set with $u$ and interacted user set with $i$, respectively. LightGCN only retains the aggregation of neighbors’ representation but removes the aggregation of interaction information representation. The final embeddings are calculated as:
(4) $\mathbf{\hat{e}}_{u}=\sum_{l=0}^{L}\frac{\mathbf{e}_{u}^{(l)}}{L+1},\quad\mathbf{\hat{e}}_{i}=\sum_{l=0}^{L}\frac{\mathbf{e}_{i}^{(l)}}{L+1},$
where $L$ is the total layer number. LightGCN considers the ego layer-00 cause it removes the first term $\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}$ and $\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}$ in NGCF.
We argue that observation 4 is the main reason why LightGCN performs better than NGCF in most cases. Observations 1-3 were not sufficient to verify that the entire feature transformation and nonlinear operation do not contribute to feature extraction. Besides, we argue that LightGCN verified the feature transformation not contributing to learning better features by comparing NGCF with removed both $\mathbf{W}_{1}$ and the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$, which is not a fair comparison. We point out that $\mathbf{e}_{u}^{(l)}$ and $\mathbf{e}_{i}^{(l)}$ naturally contain information about user preferences and item properties that can be adequately described through multiple feature dimensions. Therefore, $\mathbf{W}_{1}$ is an unnecessary feature transformation part. However, we point out that the interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ contains valuable interaction information, which can not be easily extracted by the heuristic rule. In this case, $\mathbf{W}_{2}$ is a necessary feature transformation part that contributes to feature extraction. In the next subsection, we provide empirical evidence for these arguments.
Table 1. Performance comparison of LightGCN, NGCF, and four variants of NGCF in terms of Recall@20 and NDCG@20.
Dataset Metrics LightGCN NGCF NGCF-f1 NGCF-f2 NGCF-i NGCF-n
MOOC Recall@20 0.3307 0.3361 0.3377$\uparrow$ 0.3357$\downarrow$ 0.3301$\downarrow$ 0.3343$\downarrow$
NDCG@20 0.1811 0.1894 0.1926$\uparrow$ 0.1897$\uparrow$ 0.1824$\downarrow$ 0.1878$\downarrow$
Games Recall@20 0.0447 0.0379 0.0414$\uparrow$ 0.0388$\uparrow$ 0.0362$\downarrow$ 0.0373$\downarrow$
NDCG@20 0.0227 0.0196 0.0209$\uparrow$ 0.0192$\downarrow$ 0.0181$\downarrow$ 0.0196--
@@ 2.3. Re-Analysis for NGCF
We implement four simplified variants of NGCF:
  • NGCF-f1, which removes feature transformation matrix $\mathbf{W}_{1}$.
  • NGCF-f2, which removes feature transformation matrix $\mathbf{W}_{2}$.
  • NGCF-i, which removes the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$.
  • NGCF-n, which removes the nonlinear operation $\sigma$.
We keep all hyper-parameter settings the same as the optimal settings of NGCF. As Table 1 shows, we conclude the findings that:
1. For both MOOC and Games datasets, removing the feature transformation matrix $\mathbf{W}_{1}$ will lead to observed improvements. Therefore, the feature transformation matrix $\mathbf{W}_{1}$ is unnecessary.
2. For MOOC dataset, removing the feature transformation matrix $\mathbf{W}_{2}$ leads to a slight deterioration. However, for Games dataset, it will lead to a small improvement. We own this to that user-item interaction in MOOC dataset has higher density than in Games dataset.
3. Removing the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ will lead to a performance degradation on all datasets. It shows that this part can bring valuable interaction information.
4. Nonlinear operation only brings a small positive effect.
5. NGCF outperforms LightGCN on the MOOC dataset verifying that NGCF has better representation power than LightGCN, but the training is more difficult.
Then, we can draw some conclusions. First, the feature transformation matrix $\mathbf{W}_{1}$ is unnecessary for NGCF. Besides, the feature transformation matrix $\mathbf{W}_{2}$ and the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ is beneficial for feature extraction. Last, nonlinear operation occupies only a minor influence.
@@ 3. Methodology
In this section, we detail the overall architecture of our FourierKAN-GCF, NGCF, and LightGCN in Figure 1. Our Fourier Kolmogorov-Arnold Network provides a powerful interaction representation aggregation part in NGCF to achieve a significant performance improvement and make it easy to train.
Figure 1. Overview of FourierKAN-GCF, NGCF, and LightGCN
@@ 3.1. Kolmogorov-Arnold Network (KAN)
Kolmogorov-Arnold Network (Liu et al., 2024) is a promising alternative to Multi-Layer Perceptron (MLP). MLP is inspired by the universal approximation theorem (Cybenko, 1989). KAN focuses on the Kolmogorov-Arnold representation theorem (Kolmogorov, 1961). KAN can be formed as:
@@
where $\phi_{q,p}$ are univariate functions that map each input variable $\mathbf{x}_{p}$ such $\phi_{q,p}:[0,1]\rightarrow\mathbb{R}$ and $\phi_{q}:\mathbb{R}\rightarrow\mathbb{R}$. $\phi_{q,p}(\mathbf{x}_{p})$ is trainable activation function. In the KolmogovArnold theorem, the inner functions form a KAN layer with $n_{in}$ = $n$ and $n_{out}$ = $12n+$ 1, and the outer functions form a KAN layer with $n_{in}$ = $12n+$ 1 and $n_{out}$ = $n$. So the Kolmogorov-Arnold representations in this formula are simply compositions of two KAN layers. A useful trick is that it includes a basis function $b(\mathbf{x})$ such that the activation function $\phi(\mathbf{x})$ is the sum of the basis function $b(\mathbf{x})$ and the spline function:
(6) $\phi(\mathbf{x})=\mathbf{w}(b(\mathbf{x})+\operatorname{spline}(\mathbf{x})).$
$\operatorname{spline}(\mathbf{x})$ is parameterized as a linear combination of B-splines:
(7) $b(x)=\operatorname{silu}(\mathbf{x})=\frac{x}{1+e^{-\mathbf{x}}},\quad\operatorname{spline}(\mathbf{x})=\sum_{i}c_{i}B_{i}(\mathbf{x}),$
where $c_{i}$ is trainable.

### v3

# Enhancing Graph Collaborative Filtering with FourierKAN Feature Transformation

@@ Abstract.
Graph Collaborative Filtering (GCF) has emerged as a dominant paradigm in modern recommendation systems, excelling at modeling complex user-item interactions and capturing high-order collaborative signals through graph-structured learning. Most existing GCF models predominantly rely on simplified graph architectures like LightGCN, which strategically remove feature transformation and activation functions from vanilla graph convolution networks. Through systematic analysis, we reveal that feature transformation in message propagation can enhance model representation, though at the cost of increased training difficulty. To this end, we propose FourierKAN-GCF, a novel GCN framework that adopts Fourier Kolmogorov-Arnold Networks as efficient transformation modules within graph propagation layers. This design enhances model representation while decreasing training difficulty. Our FourierKAN-GCF can achieve higher recommendation performance than most widely used GCF backbone models. In addition, it can be integrated into existing advanced self-supervised models as a backbone, replacing their original backbone to achieve enhanced performance. Extensive experiments on three public datasets demonstrate the superiority of FourierKAN-GCF.
@@
journalyear: 2025copyright: acmlicensedconference: Proceedings of the 34th ACM International Conference on Information and Knowledge Management; November 10–14, 2025; Seoul, Republic of Koreabooktitle: Proceedings of the 34th ACM International Conference on Information and Knowledge Management (CIKM ’25), November 10–14, 2025, Seoul, Republic of Koreadoi: 10.1145/3746252.3760909isbn: 979-8-4007-2040-6/2025/11
@@ 1. Introduction
Recommender systems are widely used to alleviate information overload on the Web (Chen et al., 2020; He et al., 2017; Chen et al., 2025a, b; Xu et al., 2025e, a), aiming to recommend suitable items for users based on their historical behavior. Collaborative filtering (CF) addresses this by learning user preferences from similar users. Recently, graph-based CF (GCF) models (He et al., 2020; Wang et al., 2019; Jiang et al., 2023; Xu et al., 2025d, c, b, 2024b) have achieved notable success, as user-item interactions are naturally graph-structured. For example, NGCF (Wang et al., 2019) adopts whole standard GCN in recommender systems, which retains the feature transformation and nonlinear operation. However, LightGCN (He et al., 2020) states that both feature transformation and nonlinear operation are unnecessary in the recommendation field and further proposes a lightweight GCN. Consequently, most subsequent GCF models (Chen et al., 2020; Cai et al., 2023; Xu et al., 2024a) adopt LightGCN as the backbone for further exploration.
@@
We provide an empirical analysis for NGCF and LightGCN in Section 2. Then, we state that feature transformation in NGCF can enhance the interaction representation and boost the performance of GCN, but increases the training difficulty. In this work, we introduce a simple yet powerful graph-based recommendation model called FourierKAN-GCF. Specifically, FourierKAN-GCF incorporates a unique Fourier Kolmogorov-Arnold Network (KAN) in place of the traditional multilayer perceptron (MLP) within the feature transformation during message passing in GCNs. This substitution enhances the representational capabilities and reduces the difficulty of training for GCFs. FourierKAN-GCF can achieve higher recommendation performance than most widely-used GCF backbone models. In addition, FourierKAN-GCF can be integrated into existing advanced self-supervised models as a backbone, replacing their original backbone to achieve enhanced performance. Extensive experiments on public datasets demonstrate the superiority of FourierKAN-GCF over state-of-the-art methods. Our work is intended to reawaken researchers’ thinking about feature transformation in GCF, rather than arbitrarily removing it.
@@ 2.1. NGCF Brief
NGCF (Wang et al., 2019) retains feature transformation and nonlinear operation during the message passing in GCN, formally:
@@
where $\mathbf{e}_{u}^{(l)}$ and $\mathbf{e}_{i}^{(l)}$ represent the embedding of user $u$ and item $i$ after $l$ layers message propagation, respectively. $\mathcal{N}_{u}$ and $\mathcal{N}_{i}$ denote the interacted item set with $u$ and interacted user set with $i$, respectively. $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ are two trainable weight matrices to perform feature transformation in each layer. $\sigma$ is the nonlinear activation function. It is worth noting that $\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}$ and $\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}$ can be regarded as the aggregated representation from neighbors, while $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ and $\mathbf{W}_{2}(\mathbf{e}_{u}^{(l)}\odot\mathbf{e}_{i}^{(l)})$ can be regarded as the aggregated representation from interaction information. The final embeddings are calculated by $\mathbf{\hat{e}}_{u}=\{\mathbf{e}_{u}^{(1)}||\mathbf{e}_{u}^{(2)}||...||\mathbf{e}_{u}^{(L)}\}$ and $\mathbf{\hat{e}}_{i}=\{\mathbf{e}_{i}^{(1)}||\mathbf{e}_{i}^{(2)}||...||\mathbf{e}_{0}^{(L)}\}$, where $||$ is the concatenation operation. NGCF only concatenates layer-$1$ to layer-$L$ and ignores ego layer-0, since ego layer-0 has already been considered in the first term $\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}$ and $\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}$ of message passing and propagation.
@@ 2.2. LightGCN Brief
LightGCN (He et al., 2020) analyzes the feature transformation and nonlinear operation in NGCF. It offers four observations: a) Removing the entire feature transformation, including $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$, leads to consistent improvements over NGCF. b) Removing only nonlinear operation $\sigma$ will lead to a small deterioration of performance. c) Removing both entire feature transformation and nonlinear operation can improve performance significantly. d) The deterioration of NGCF stems from the training difficulty rather than over-fitting.
To this end, LightGCN removes feature transformation and nonlinear operation during the message passing in GCN. Formally, the user-item graph to propagate embeddings as:
@@
where $\mathbf{e}_{u}^{(l)}$ and $\mathbf{e}_{i}^{(l)}$ represent the embedding of user $u$ and item $i$ after $l$ layers message propagation, respectively. $\mathcal{N}_{u}$ and $\mathcal{N}_{i}$ denote the interacted item set with $u$ and interacted user set with $i$, respectively. LightGCN only retains the aggregation of neighbors’ representation but removes the aggregation of interaction information representation. The final embeddings are calculated as $\mathbf{\hat{e}}_{u}=\sum_{l=0}^{L}\frac{1}{L+1}\mathbf{e}_{u}^{(l)}$ and $\mathbf{\hat{e}}_{i}=\sum_{l=0}^{L}\frac{1}{L+1}\mathbf{e}_{i}^{(l)}$, where $L$ is the total layer number. LightGCN considers the ego layer-0 cause it removes the first term $\mathbf{W}_{1}\mathbf{e}_{u}^{(l)}$ and $\mathbf{W}_{1}\mathbf{e}_{i}^{(l)}$.
We argue that observation (d) is the main reason why LightGCN performs better than NGCF in most cases. Observations (a-c) were not sufficient to verify that the entire feature transformation and nonlinear operation do not contribute to feature extraction. Besides, we argue that LightGCN verified the feature transformation not contributing to learning better features by comparing NGCF with removed both $\mathbf{W}_{1}$ and the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$, which is not a fair comparison. We point out that $\mathbf{e}_{u}^{(l)}$ and $\mathbf{e}_{i}^{(l)}$ naturally contain information about user preferences and item properties that can be adequately described through multiple feature dimensions. Therefore, $\mathbf{W}_{1}$ is an unnecessary feature transformation part. However, we point out that the interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ contains valuable interaction information, which can not be easily extracted by the heuristic rule. In this case, $\mathbf{W}_{2}$ is a necessary feature transformation part that contributes to feature extraction. In the next subsection, we provide an empirical analysis.
Table 1. Performance comparison of LightGCN, NGCF, and six variants of NGCF in terms of Recall@20 (R@20) and NDCG@20 (N@20).
Datasets MOOC Amazon Gowalla
Metrics R@20 N@20 R@20 N@20 R@20 N@20
LightGCN 0.3307 0.1811 0.0447 0.0227 0.1830 0.1152
NGCF 0.3361 0.1894 0.0379 0.0196 0.1755 0.1013
[…]
@@ 2.3. Re-Analysis for NGCF
We implement six variants of NGCF: 1) NGCF-f1 removes feature transformation matrix $\mathbf{W}_{1}$. 2) NGCF-f2 removes feature transformation matrix $\mathbf{W}_{2}$. 3) NGCF-i removes the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$. 4) NGCF-f1-f2 removes both $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$. 5) NGCF-f1-i removes both $\mathbf{W}_{1}$ and $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$. 6) NGCF-n removes the nonlinear operation $\sigma$.
We keep all optimal hyper-parameter settings as the NGCF reported. As Table 1 shows, we conclude the findings: a) For all three datasets, removing the feature transformation matrix $\mathbf{W}_{1}$ will lead to observed improvements. Therefore, the feature transformation matrix $\mathbf{W}_{1}$ is unnecessary. b) For the MOOC dataset, removing the feature transformation matrix $\mathbf{W}_{2}$ leads to a slight deterioration. However, for the other two datasets, it will lead to a small improvement. Moreover, removing both $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ will lead to a small deterioration than only removing $\mathbf{W}_{1}$ for all datasets. We own this phenomenon to that removing only $\mathbf{W}_{1}$ can make the model optimization focus on training $\mathbf{W}_{2}$, and remove both $\mathbf{W}_{1}$ and $\mathbf{W}_{2}$ will lose the representation power from feature transformation. c) Removing the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ will lead to a performance degradation on all datasets. Additionally, removing both $\mathbf{W}_{1}$ and $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ will make a obviously degration than only remove $\mathbf{W}_{1}$ and even NGCF. This shows that this part can bring valuable information about interaction. d) Nonlinear operation only brings a small positive effect. e) NGCF outperforms LightGCN on the MOOC dataset verifying that NGCF has better representation power than LightGCN, but the training is more difficult (He et al., 2020). Then, we can draw some conclusions. First, the feature transformation matrix $\mathbf{W}_{1}$ is unnecessary for NGCF. Besides, the feature transformation matrix $\mathbf{W}_{2}$ and the whole interaction representation aggregation part $\mathbf{W}_{2}(\mathbf{e}_{i}^{(l)}\odot\mathbf{e}_{u}^{(l)})$ is beneficial for feature extraction. Last, nonlinear operation only has a minor influence.
To this end, a feature transformation approach, simpler to train than MLPs but with strong representational capabilities, could be effective in graph learning for recommendations. KAN (Liu et al., 2024) is regarded as a promising alternative to MLP. While KAN shares the same theoretically unlimited representational capacity as MLP (Wang et al., 2024), the practical representational capacity of MLP is constrained by the hidden dimensionality. In contrast, the practical representational capacity of KAN depends on its ability to fit trainable activation functions. Our work can be seen as exploring adopting KAN as a feature transformation approach within GCF.
@@ 3. Methodology
We detail the overview of FourierKAN-GCF [^1] Code is available at: https://anonymous.4open.science/r/FourierKAN-GCF-r. in Figure 1. FourierKAN-GCF introduces a promising feature transformation for GCF, which significantly boosts performance and simplifies training. Note that FourierKAN-GCF can be adopted as the backbone in existing advanced self-supervised models.
Figure 1. Overview of FourierKAN-GCF.
@@ 3.1. Kolmogorov-Arnold Network (KAN)
Kolmogorov-Arnold Network (Liu et al., 2024) is a promising alternative to Multi-Layer Perceptron (MLP). MLP is inspired by the universal approximation theorem (Cybenko, 1989). KAN focuses on the Kolmogorov-Arnold representation theorem (Kolmogorov, 1961). Specifically, unlike MLPs, which have fixed activation functions on nodes, KANs contain learnable activation functions on edges (weights). This unique architecture enables KANs to learn nonlinear functions more effectively. Formally:
@@
where $\phi_{q,p}$ are univariate functions that map each input variable $\mathbf{x}_{p}$ such $\phi_{q,p}:[0,1]\rightarrow\mathbb{R}$ and $\phi_{q}:\mathbb{R}\rightarrow\mathbb{R}$. $\phi_{q,p}(\mathbf{x}_{p})$ is trainable activation function. In the KolmogovArnold theorem, the inner functions form a KAN layer with $_{in} = nn$ and $n_{out}$ = $2n+1$, and the outer functions form a KAN layer with $n_{in}$ = $2n+1$ and $_{out} = nn$. So the Kolmogorov-Arnold representations in this formula are simply compositions of two KAN layers. A useful trick is that it includes a basis function $b(\mathbf{x})$ such that the activation function $\operatorname{silu}(\mathbf{x})=\frac{x}{1+e^{-\mathbf{x}}}$. $\phi(\mathbf{x})=\mathbf{w}(b(\mathbf{x})+\operatorname{spline}(\mathbf{x}))$ is the sum of the basis function $b(\mathbf{x})$ and function $\operatorname{spline}(\mathbf{x})=\sum_{i}c_{i}B_{i}(\mathbf{x})$ is a linear combination of B-splines, where $c_{i}$ is a trainable parameter.